# MDSGen: Fast and Efficient Masked Diffusion Temporal-Aware Transformers for Open-Domain Sound Generation

### v1

@@ Abstract
We introduce MDSGen, a novel framework for vision-guided open-domain sound generation optimized for model parameter size, memory consumption, and inference speed. This framework incorporates two key innovations: (1) a redundant video feature removal module that filters out unnecessary visual information, and (2) a temporal-aware masking strategy that leverages temporal context for enhanced audio generation accuracy. In contrast to existing resource-heavy Unet-based models, MDSGen employs denoising masked diffusion transformers, facilitating efficient generation without reliance on pre-trained diffusion models. Evaluated on the benchmark VGGSound dataset, our smallest model (5M parameters) achieves $97.9$% alignment accuracy, using 172×172$\times$172 × fewer parameters, $371$% less memory, and offering 36×36$\times$36 × faster inference than the current 860M-parameter state-of-the-art model ($93.9$% accuracy). The larger model (131M parameters) reaches nearly $99$% accuracy while requiring 6.5×6.5$\times$6.5 × fewer parameters. These results highlight the scalability and effectiveness of our approach.
@@ 1 Introduction
Vision-guided audio generation has gained significant attention due to its crucial role in Foley sound synthesis for the video and film production industry (Ament, 2014). This paper focuses on Video-to-Audio (V2A) generation, a key task not only for adding realistic sound to silent videos created by emerging text-to-video models (Blattmann et al., 2023; Khachatryan et al., 2023; Huang et al., 2024; Ouyang et al., 2024) but also for enhancing practical applications in professional video production. Sound generation is essential for creating immersive experiences and achieving seamless audio-visual synchronization. However, achieving both semantic alignment and temporal synchronization in V2A remains a significant challenge. Previous approaches, such as GAN-based methods (Chen et al., 2020b) and Transformer-based autoregressive models (Iashin & Rahtu, 2021), have struggled with synchronizing audio to content while maintaining relevance. Diff-Foley (Luo et al., 2023) improved this by employing contrastive learning for video-audio alignment and leveraging diffusion models, achieving impressive sound quality. Other methods like See and Hear (Xing et al., 2024) and FoleyCrater (Zhang et al., 2024) utilize large pre-trained models for high-quality audio generation. However, these models rely on hundreds of millions of parameters. In contrast, our work demonstrates that a much smaller model can deliver high performance (see Fig. 1). Most existing approaches rely on Unet architectures, which present scalability limitations. Additionally, current methods often use video features that include redundant information. In contrast, we propose MDSGen, a novel framework for open-domain sound synthesis based on a pure Transformer architecture. MDSGen incorporates a temporal-aware masking scheme and a redundant feature removal module, enabling it to achieve superior performance while being significantly more efficient. Further analysis highlights the effectiveness of our approach with compelling evidence of its advantages. Our contributions are as follows:
@@ 2.1 Open-domain Sound Generation
Diffusion-based Approach. To overcome these limitations, Diff-Foley (Luo et al., 2023) introduced a two-stage method that enhances semantic and temporal alignment via contrastive pre-training on aligned video-audio pairs, followed by latent diffusion for improved inference efficiency. Similarly, See and Hear (Xing et al., 2024) utilizes ImageBind (Girdhar et al., 2023) and AudioLDM (Liu et al., 2023) for various audio tasks, while FoleyCrafter (Zhang et al., 2024) combines a pre-trained text-to-audio model with a ControlNet-style module (Zhang et al., 2023) for high-quality, synchronized Foley generation. Although these diffusion approaches show promise, they often rely on large models with hundreds of millions of parameters and predominantly utilize U-Net architectures, leaving the potential of transformer-based architectures largely untapped.
Our proposed method leverages diffusion transformers (Peebles & Xie, 2023) and masking techniques for efficient learning. It also addresses the issue of redundant video features in Diff-Foley (Luo et al., 2023), which hinders further improvements in audio generation.
@@ 2.2 Latent Masked Diffusion Transformers
The Denoising Diffusion Transformer (DiT) introduced by Peebles & Xie (2023) replaces the traditional U-Net with a fully transformer-based architecture for latent diffusion, demonstrating remarkable performance in large-scale image generation on ImageNet. Following this, Gao et al. (2023) proposed the Masked Diffusion Transformer (MDT), which enhances ImageNet generation through spatial context-aware masking. Inspired by MDT, Pham et al. (2024) developed X-MDPT, using cross-view masking to establish correspondence between pose and reference images for improved person image generation. Additionally, MDT-A2G (Mao et al., 2024) explored masked diffusion transformers for gesture generation, while QA-MDT (Li et al., 2024) adapted this technique for music generation.
In contrast to these works, we focus on lightweight masked diffusion models for video-guided audio generation, introducing temporal-aware masking for audio and a design that removes redundant video features to enhance generation effectiveness.
@@ 3 Method
We aim to develop a simple yet effective framework for vision-guided sound generation using transformers, addressing the limitations of existing approaches that rely on traditional U-Net architectures, which are less scalable and efficient. Our framework, illustrated in Fig. 2, consists of a novel Vision Extractor with a learnable Reducer that captures essential information from video input to generate a concise conditional output for the denoising diffusion process. This is followed by a Denoising Diffusion Transformer that maps Gaussian noise to sound distributions based on the extracted visual features. Additionally, we introduce a Masked Temporal-Aware Network (MTANet) that serves as a strong regularization mechanism for the transformer, enhancing overall performance. Finally, we incorporate Channel Selection for mel-spectrograms that further improves results.
@@ 3.1 Denoising Diffusion Transformer
We adopt the DiT backbone introduced by Peebles & Xie (2023), which leverages transformers with self-attention during the denoising diffusion training process. Given an audio signal $\mathbf{A}\in\mathbb{R}^{L_{\mathbf{A}}}$ of length $L_{\mathbf{A}}$ and a silent video $\mathbf{V}\in\mathbb{R}^{L_{V}\times 3\times 224\times 224}$ of length $L_{V}$, the audio is first transformed into a mel-spectrogram $\mathbf{X}\in\mathbb{R}^{128\times 512}$, while the video is encoded into $\mathbf{v}\in\mathbb{R}^{L_{V}\times 512}$ and further reduced to $\vec{v}\in\mathbb{R}^{1\times D}$. The mel-spectrogram is repeated across 3 channels, forming $\mathbf{X^{\prime}}\in\mathbb{R}^{3\times 128\times 512}$, and passed through the VAE from Stable Diffusion (Rombach et al., 2022; Luo et al., 2023) to obtain a latent embedding $\mathbf{x}\in\mathbb{R}^{4\times 16\times 64}$. This latent representation is patched and tokenized into image tokens using a patch size of $p=2$ (DiT’s default), resulting in $\mathbf{x}^{\prime}\in\mathbb{R}^{256\times D}$, where $L_{\mathbf{x^{\prime}}}=256$ and $D=768$ for the B-size model. These tokens are then fed into the $N=N_{1}+N_{2}$ self-attention layers of the Transformer to predict the noise $\epsilon$ added to the latent $\mathbf{x}$. Conditioned on the video encoding $\vec{v}$, the Transformer model $\phi$ learns the distribution $p_{\phi}(\mathbf{x}|\vec{v})$. During training, Gaussian noise $\epsilon\in\mathcal{N}(0,\mathbf{I})$ is added to the latent $\mathbf{x}$ to generate $\mathbf{x}_{t}$ at timestep $t\in[1,T]$. The overall training objective is:
@@
Here, $\lambda$ is the balance factor between the standard denoising diffusion loss (the first term in Eq. 1) and the masking loss (the second term), with $\lambda=1.0$ for optimal performance. The masking function $\mathcal{M}_{\phi}$, which includes the MTANet introduced later, applies temporal-aware masking. During inference, given a silent video, the model starts from Gaussian noise (no audio provided), and the predicted latent $\hat{\mathbf{x}}\in\mathbb{R}^{4\times 16\times 64}$ is iteratively denoised and decoded by the VAE decoder to recover the mel-spectrogram $\widehat{\mathbf{X}}_{RGB}\in\mathbb{R}^{3\times 128\times 512}$. Channel selection refines this into $\widehat{\mathbf{X}}\in\mathbb{R}^{128\times 512}$, and the final waveform is reconstructed from the mel-spectrogram using the Griffin-Lim algorithm (Griffin & Lim, 1984), consistent with previous work (Luo et al., 2023) for fair comparison.
@@ 3.2 Vision Extractor
The second key component of our framework is the Vision Extractor with a learnable Reducer network that aligns video features with audio while condensing temporal information. We leverage the pre-trained CAVP model from (Luo et al., 2023), which was trained on Audioset using contrastive loss to extract video features aligned with audio. However, we identified that the CAVP features contain redundancies that could negatively impact generation quality. Diff-Foley (Luo et al., 2023) linearly maps original feature dimensions from $\mathbf{v}\in\mathbb{R}^{L_{V}\times 512}$ to $\mathbf{v}\in\mathbb{R}^{L_{V}\times 768}$ and retains this full dimensionality during the latent diffusion process via cross-attention, with $L_{V}$ is the video feature length.
Our approach reduces the dimensionality to $\vec{v}\in\mathbb{R}^{1\times 768}$, offering more concise and efficient information for denoising diffusion. Specifically, we project the encoded features $L_{V}\times 512$ through a multi-layer perceptron (MLP) into the transformer feature space ($L_{V}\times 768$ for the size B-model). These features are then passed through a reducer module, an $1\times 1$ convolutional layer, which condenses the high-dimensional features into a lightweight form $\vec{v}\in\mathbb{R}^{1\times 768}$. This compact representation is integrated into the denoising diffusion process through Adaptive LayerNorm (AdaLN) modulation. Unlike Diff-Foley, which retains all $L_{V}$ video channels representing $L_{V}$ video frames for cross-attention within the Stable Diffusion U-Net backbone, our method minimizes redundant features that could lead to overfitting. Our analysis shows that the $L_{V}$-frame input features share over $90$% similarity, indicating considerable redundancies, which we further address in our ablation study.
@@ 3.3 Audio Masked Temporal-Aware Network
Figure 3: Audio Masking Strategies. Here, the red square is the learnable mask token.
Thirdly, we introduce a novel technique that exploits the sound information’s natural characteristic: the temporal sense. The existing masking, Spatial-Aware Mask (SAM) proposed by MDT (Gao et al., 2023) is designed for image data to learn the spatial context within the image. But here, in the audio data (represented by mel-spectrogram with 2D data), the SAM masking method yields a sub-optimal solution because it cannot model the exact nature of temporal meaning in the audio data.
To overcome this limitation, we propose the Temporal-Aware Mask (TAM) strategy instead of SAM, which tries to mask the whole set of tokens along the temporal dimension. As shown in the ablation section, this novel masking helps significantly boost performance in all metrics compared to the existing method SAM specified for image data as shown in Fig. 3. Interestingly, despite this simple strategy, it can help the denoising transformer models learn to generate audio much better than random masks as used in existing MDT designed for image data.
@@ 3.4 Classifier-Free Guidance
3.5 Mel-Spectrogram Channel Selection
We discovered that the audio extracted from the RGB output channels of the VAE decoder exhibits different qualities. Since the mel-spectrogram is 2D, we treat it as a single gray channel and duplicate it into three channels before feeding it into the VAE, which is pre-trained to reconstruct RGB images. At the decoding stage, the VAE outputs three channels: $\widehat{\mathbf{X}}_{RGB}\in\mathbb{R}^{3\times 128\times 512}$, with $\widehat{\mathbf{X}}_{RGB}[i,:,:]\in\mathbb{R}^{128\times 512},i\in\{0,1,2\}$ representing the R, G, and B channels. While Diff-Foley (Luo et al., 2023) used the first channel (R) as their final output, our empirical tests reveal that the second channel (G) consistently performs better. We further ablated channel selection and found that using the G channel yields superior results compared to the R channel used in prior work.
@@ 4.1 Dataset and Evaluation Metrics
(i) Dataset. We evaluate our method on the VGGSound dataset (Chen et al., 2020a), using the original train/test splits with 175k and 15k samples, respectively, and on the Flick-SoundNet dataset Aytar et al. (2016) with 5k test samples. (ii) Metrics. We first use the same metrics as prior work (Luo et al., 2023), including FID, IS, KL, and Alignment Accuracy, using their provided scripts for Align. Acc. and SpecVQGAN code for FID, IS, and KL. Second, we assess general vision-audio alignment in the Image2Audio task using CIoU and AUC metrics with scripts from (Mo & Morgado, 2022). For this, we generate pseudo videos by replicating a single input image for all frames. Third, we compare efficiency using parameter count, memory usage, and inference speed.

### v2

@@ Abstract
We introduce MDSGen, a novel framework for vision-guided open-domain sound generation optimized for model parameter size, memory consumption, and inference speed. This framework incorporates two key innovations: (1) a redundant video feature removal module that filters out unnecessary visual information, and (2) a temporal-aware masking strategy that leverages temporal context for enhanced audio generation accuracy. In contrast to existing resource-heavy Unet-based models, MDSGen employs denoising masked diffusion transformers, facilitating efficient generation without reliance on pre-trained diffusion models. Evaluated on the benchmark VGGSound dataset, our smallest model (5M parameters) achieves $97.9$% alignment accuracy, using 172×172$\times$172 × fewer parameters, $371$% less memory, and offering 36×36$\times$36 × faster inference than the current 860M-parameter state-of-the-art model ($93.9$% accuracy). The larger model (131M parameters) reaches nearly $99$% accuracy while requiring 6.5×6.5$\times$6.5 × fewer parameters. These results highlight the scalability and effectiveness of our approach. The code is available at https://bit.ly/mdsgen.
@@ 1 Introduction
Vision-guided audio generation has gained significant attention due to its crucial role in Foley sound synthesis for the video and film production industry (Ament, 2014). This paper focuses on Video-to-Audio (V2A) generation, a key task not only for adding realistic sound to silent videos created by emerging text-to-video models (Blattmann et al., 2023; Khachatryan et al., 2023; Huang et al., 2024; Ouyang et al., 2024) but also for enhancing practical applications in professional video production. Sound generation is essential for creating immersive experiences and achieving seamless audio-visual synchronization. However, achieving both semantic alignment and temporal synchronization in V2A remains a significant challenge. Previous approaches, such as GAN-based methods (Chen et al., 2020b) and Transformer-based autoregressive models (Iashin & Rahtu, 2021), have struggled with synchronizing audio to content while maintaining relevance. Diff-Foley (Luo et al., 2023) improved this by employing contrastive learning for video-audio alignment and leveraging diffusion models, achieving impressive sound quality. Other methods like See and Hear (Xing et al., 2024) and FoleyCrater (Zhang et al., 2024) utilize large pre-trained models for high-quality audio generation. However, these models rely on hundreds of millions of parameters. In contrast, our work demonstrates that a much smaller model can deliver high performance (see Fig. 1). Most existing approaches rely on Unet architectures, which present scalability limitations. Additionally, current methods often use video features that include redundant information.
In contrast, we propose MDSGen, a novel framework for open-domain sound synthesis based on a pure Transformer architecture. MDSGen incorporates a temporal-aware masking scheme and a redundant feature removal module, enabling it to achieve superior performance while being significantly more efficient. Further analysis highlights the effectiveness of our approach with compelling evidence of its advantages. Our contributions are as follows:
@@ 2.1 Open-domain Sound Generation
Diffusion-based Approach. To overcome these limitations, Diff-Foley (Luo et al., 2023) introduced a two-stage method that enhances semantic and temporal alignment via contrastive pre-training on aligned video-audio pairs, followed by latent diffusion for improved inference efficiency. Similarly, See and Hear (Xing et al., 2024) utilizes ImageBind (Girdhar et al., 2023) and AudioLDM (Liu et al., 2023) for various audio tasks, while FoleyCrafter (Zhang et al., 2024) combines a pre-trained text-to-audio model with a ControlNet-style module (Zhang et al., 2023) for high-quality, synchronized Foley generation. Although these diffusion approaches show promise, they often rely on large models with hundreds of millions of parameters and predominantly utilize U-Net architectures, leaving the potential of transformer-based architectures largely untapped. Our proposed method leverages diffusion transformers (Peebles & Xie, 2023) and masking techniques for efficient learning. It also addresses the issue of redundant video features in Diff-Foley (Luo et al., 2023), which hinders further improvements in audio generation.
@@ 2.2 Latent Masked Diffusion Transformers
The Denoising Diffusion Transformer (DiT) introduced by Peebles & Xie (2023) replaces the traditional U-Net with a fully transformer-based architecture for latent diffusion, demonstrating remarkable performance in large-scale image generation on ImageNet. Following this, Gao et al. (2023) proposed the Masked Diffusion Transformer (MDT), which enhances ImageNet generation through spatial context-aware masking. Inspired by MDT, Pham et al. (2024) developed X-MDPT, using cross-view masking to establish correspondence between pose and reference images for improved person image generation. Additionally, MDT-A2G (Mao et al., 2024) explored masked diffusion transformers for gesture generation, while QA-MDT (Li et al., 2024) adapted this technique for music generation. In contrast to these works, we focus on lightweight masked diffusion models for video-guided audio generation, introducing temporal-aware masking for audio and a design that removes redundant video features to enhance generation effectiveness.
2.3 Masked Data in the Audio Modality
Several works have explored masking techniques for audio processing. MaskVAT (Pascual et al., 2024) introduces a V2A system that integrates a full-band audio codec, masked generative modeling, and multi-modal features to enhance audio quality, semantic alignment, and temporal synchronization. SoundStorm (Borsos et al., 2023) employs the non-autoregressive MaskGIT (Chang et al., 2022) approach for efficient text-to-audio generation. Similarly, VamNet (Garcia et al., 2023) applies MaskGIT to music generation, while (Bai et al., 2022) uses masking in the pixel space of mel-spectrograms for non-diffusion-based text-to-audio tasks. These approaches differ fundamentally from our diffusion-based framework, which applies masking in the latent space with added Gaussian noise. AudioMAE (Huang et al., 2022) and SpecAugment (Park et al., 2019) share similarities with our method in employing masking for audio data. However, key distinctions exist: both focus on masking in the pixel space of clean mel-spectrograms for representation learning in downstream recognition tasks. In contrast, our approach utilizes masking in the latent space of a VAE within the diffusion framework, targeting audio generation.
@@ 3 Method
We aim to develop a simple yet effective framework for vision-guided sound generation using transformers, addressing the limitations of existing approaches that rely on traditional U-Net architectures, which are less scalable and efficient. Our framework, illustrated in Fig. 2, consists of a novel Vision Extractor with a learnable Reducer that captures essential information from video input to generate a concise conditional output for the denoising diffusion process. Next, a Denoising Diffusion Transformer maps Gaussian noise to sound distributions using extracted visual features. We also introduce a Masked Temporal-Aware Network (MTANet) for regularization, boosting performance. Finally, channel selection for mel-spectrograms, which enhances results with image VAEs, is optional.
@@ 3.1 Denoising Diffusion Transformer
Our method supports both audio- and image-based VAEs. For instance, we describe using an image VAE (Luo et al., 2023; Chen et al., 2024), and for audio VAEs like AudioLDM, we adjust the three channels to one. We adopt the DiT backbone introduced by Peebles & Xie (2023) for denoising diffusion training. Given an audio signal $\mathbf{A}\in\mathbb{R}^{L_{\mathbf{A}}}$ of length $L_{\mathbf{A}}$ and a silent video $\mathbf{V}\in\mathbb{R}^{L_{V}\times 3\times 224\times 224}$ of length $L_{V}$, the audio is first transformed into a mel-spectrogram $\mathbf{X}\in\mathbb{R}^{128\times 512}$, while the video is encoded into $\mathbf{v}\in\mathbb{R}^{L_{V}\times 512}$ and further reduced to $\vec{v}\in\mathbb{R}^{1\times D}$. The mel-spectrogram is repeated across 3 channels, forming $\mathbf{X^{\prime}}\in\mathbb{R}^{3\times 128\times 512}$, and passed through the VAE from Stable Diffusion (Rombach et al., 2022; Luo et al., 2023) to obtain a latent embedding $\mathbf{x}\in\mathbb{R}^{4\times 16\times 64}$. This latent representation is patched and tokenized into image tokens using a patch size of $p=2$ (DiT’s default), resulting in $\mathbf{x}^{\prime}\in\mathbb{R}^{256\times D}$, where $L_{\mathbf{x^{\prime}}}=256$ and $D=768$ for the B-size model. These tokens are then fed into the $N=N_{1}+N_{2}$ self-attention layers of the Transformer to predict the noise $\epsilon$ added to the latent $\mathbf{x}$. Conditioned on the video encoding $\vec{v}$, the Transformer model $\phi$ learns the distribution $p_{\phi}(\mathbf{x}|\vec{v})$. During training, Gaussian noise $\epsilon\in\mathcal{N}(0,\mathbf{I})$ is added to the latent $\mathbf{x}$ to generate $\mathbf{x}_{t}$ at timestep $t\in[1,T]$. The overall training objective is:
@@
Here, $\lambda$ is the balance factor between the standard denoising diffusion loss (the first term in Eq. 1) and the masking loss (the second term), with $\lambda=1.0$ for optimal performance. The masking function $\mathcal{M}_{\phi}$, which includes the MTANet introduced later, applies temporal-aware masking. During inference, given a silent video, the model starts from Gaussian noise (no audio provided), and the predicted latent $\hat{\mathbf{x}}\in\mathbb{R}^{4\times 16\times 64}$ is iteratively denoised and decoded by the VAE decoder to recover the mel-spectrogram $\widehat{\mathbf{X}}_{RGB}\in\mathbb{R}^{3\times 128\times 512}$. Channel selection refines this into $\widehat{\mathbf{X}}\in\mathbb{R}^{128\times 512}$, and the final waveform is reconstructed from the mel-spectrogram using the Griffin-Lim (Griffin & Lim, 1984), or neural Hifi-GAN vocoder Kong et al. (2020).
@@ 3.2 Vision Extractor
The second key component of our framework is the Vision Extractor with a learnable Reducer network that aligns video features with audio while condensing temporal information. We leverage the pre-trained CAVP model from (Luo et al., 2023), which was trained on Audioset using contrastive loss to extract video features aligned with audio. However, we identified that the CAVP features contain redundancies that could negatively impact generation quality. Diff-Foley (Luo et al., 2023) linearly maps original feature dimensions from $\mathbf{v}\in\mathbb{R}^{L_{V}\times 512}$ to $\mathbf{v}\in\mathbb{R}^{L_{V}\times 768}$ and retains this full dimensionality during the diffusion process via cross-attention, with $L_{V}$ is the video feature length. Our approach reduces the dimensionality to $\vec{v}\in\mathbb{R}^{1\times 768}$, offering more concise and efficient information for denoising diffusion. Specifically, we project the encoded features $L_{V}\times 512$ through a multi-layer perceptron (MLP) into the transformer feature space ($L_{V}\times 768$ for the size B-model). These features are then passed through a reducer module, an $1\times 1$ convolutional layer, which condenses the high-dimensional features into a lightweight form $\vec{v}\in\mathbb{R}^{1\times 768}$. This compact representation is integrated into the denoising diffusion process through Adaptive LayerNorm (AdaLN) modulation. Our method minimizes redundant features that could lead to overfitting, as shown in the train/test alignment accuracy gap Appendix Sec. A.2. Our analysis shows that the $L_{V}$-frame input features share over $90$% similarity, indicating considerable redundancies.
@@
Figure 3: Audio Masking Strategies. Here, the red square red-square is the learnable mask token.
@@ 3.3 Audio Masked Temporal-Aware Network
Thirdly, we introduce a novel technique that exploits the sound information’s natural characteristic: the temporal sense. The existing masking, Spatial-Aware Mask (SAM) proposed by MDT (Gao et al., 2023) is designed for image data to learn the spatial context within the image. But here, in the audio data (represented by mel-spectrogram with 2D data), the SAM masking method yields a sub-optimal solution because it cannot model the exact nature of temporal meaning in the audio data. To overcome this limitation, we propose the Temporal-Aware Mask (TAM) strategy instead of SAM, which tries to mask the whole set of tokens along the temporal dimension. As shown in the ablation section, this novel masking helps significantly boost performance in all metrics compared to the existing method SAM specified for image data as shown in Fig. 3. Interestingly, despite this simple strategy, it can help the denoising transformer models learn to generate audio much better than random masks as used in existing MDT designed for image data.
@@ 3.4 Classifier-Free Guidance
3.5 VAE Choice for Mel-spectrogram
Our method supports both Audio- and Image-trained VAEs. Ablation finds that audio quality varies across the RGB output channels of image-trained VAEs. Since the mel-spectrogram is 2D, we duplicate it into three channels for Stable Diffusion VAE. At the decoding stage, the VAE outputs three channels: $\widehat{\mathbf{X}}_{RGB}\in\mathbb{R}^{3\times 128\times 512}$, with $\widehat{\mathbf{X}}_{RGB}[i,:,:]\in\mathbb{R}^{128\times 512},i\in\{0,1,2\}$ representing the R, G, and B channels. Diff-Foley (Luo et al., 2023) used the R channel as output, but our empirical tests consistently show that the G channel performs better. However, when using the audio VAE, i.e. AudioLDM VAE, channel selection is no longer required.
@@ 4.1 Dataset and Evaluation Metrics
(i) Dataset. We evaluate our method on the VGGSound dataset (Chen et al., 2020a), using the original train/test splits with 175k and 15k samples, respectively, and on the Flick-SoundNet dataset Aytar et al. (2016) with 5k test samples. (ii) Metrics. We first use the same metrics as prior work (Luo et al., 2023), including FID, IS, KL, and Alignment Accuracy, using their provided scripts for Align. Acc. and SpecVQGAN code for FID, IS, and KL. Second, we assess general vision-audio alignment in the Image2Audio task using CIoU and AUC metrics with scripts from (Mo & Morgado, 2022). Third, we compare efficiency using parameter count, memory usage, and inference speed. Lastly, we provide the FAD scores and MOS results from human evaluations in the Appendix A.